OpenStudy (anonymous):

Need help multiplying this polynomial I will post the question in a comment in a second.

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 }b +\frac{ 3 }{ 4 } )(2b - \frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

And please explain how you got the answer thanks :)

OpenStudy (anonymous):

Meant to say 2 bionomeials and not polynomials sorry ;o

OpenStudy (anonymous):